OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3/2 x + 4 and passes through the point (−6, 3) help please!!! i really don't know what im doing

OpenStudy (zzr0ck3r):

y = mx+b y = 3 x = -6 m = -2/3 b = ?

OpenStudy (anonymous):

thank you so much i appreciate it